OpenStudy (anonymous):

what is the sum of these mixed numbers ? 2 3/4 + 3 3/4 =

OpenStudy (anonymous):

6 and 2/4 or 6 and 1/2

OpenStudy (anonymous):

6 1/2

OpenStudy (anonymous):

2 + 3 = 5 3/4 + 3/4 = 6/4 = 1 1/2 5 + 1 1/2 = 6 1/2

OpenStudy (anonymous):

13/2